# DIRECT: Deep Active Learning under Imbalance and Label Noise

### 14 Dec 2023 v1

@@ Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. In this work, we propose a novel algorithm that first identifies the class separation threshold and then annotate the most uncertain examples from the minority classes, close to the separation threshold. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise. Compared to existing algorithms, our algorithm saves more than 15% of the annotation budget compared to state-of-art active learning algorithm and more than 90% of annotation budget compared to random sampling.
@@ 1 Introduction
Large-scale deep learning models is playing an increasing role across many industries. Human feedback and annotations have played a significant role in developing such systems. Increasingly, we believe the role of human would shift to annotating rare yet important cases such as safety risks. In this paper, we propose an active learning algorithm to address the class imbalance problem, where we sequentially and adaptively choose examples for annotation. Our proposed algorithm optimizes for the class-balancedness and informativeness of the annotation collected.
Inspired by previous work by Zhang et al. (2022), we propose a novel reduction of the imbalanced classification problem into an one-dimensional active learning problem by arranging unlabeled examples into a list ordered by uncertainty scores. With this reduction, we apply classic active learning literature in finding the optimal separation threshold between the minority and majority classes. We then annotate examples near this threshold to gather uncertain examples.
On the other hand, existing state-of-art algorithm like GALAXY (Zhang et al., 2022) does not allow parallel annotation, and performs poorly when label noise is high. With our novel one-dimensional reduction, we inherit the noise tolerance and batch labeling advantages of classic active learning algorithms, allowing our deep active learning algorithm DIRECT to overcome such issues.
@@
  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification.
  • Our novel algorithm addresses the prevalent imbalancedness issue by annotating a more class-balanced and informative set.
  • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), our algorithm allows parallel annotation and is tolerant to label noise. It also saves $15\%$ more annotation cost compared to GALAXY, and $90\%$ of the annotation cost compared to random sampling.
@@ 2 Related Work
Deep Active Learning Under Class Balance Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009). Many uncertainty methods active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a). More sophisticated methods have also been proposed to optimize the chosen unlabeled examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as under label noises (Khosla et al., 2022) or under class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2023a).
@@ 3.1 Notations
We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation. We let $\mathcal{X}$ denote the space of possible examples. Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown. Furthermore, we study the multi-class classification problem, so the space of labels $\mathcal{Y}:=[K]$, where each element denotes the class index and $K$ is the total number of classes. Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ in each class. We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k\in[K]}N_{k}}$.
A deep active learning algorithm iteratively chooses batches of examples for annotation. During the $t$-th iteration, the algorithm is given labeled and unlabeled sets of examples, $L_{t}$ and $U_{t}$ respectively, where $L_{t}\cup U_{t}=X$ and $L_{t}\cap U_{t}=\emptyset$. The algorithm then chooses $B$ examples from the unlabeled set $X^{(t)}\subseteq U_{t}$ for their corresponding labels $Y^{(t)}$ to be annotated. The labeled and unlabeled sets are then updated, i.e., $L_{t+1}\leftarrow L_{t}\cup X^{(t)}$ and $U_{t+1}\leftarrow U_{t}\backslash X^{(t)}$. Based on new labeled set $L_{t+1}$ and its corresponding labels, a neural network $f_{t}:\mathcal{X}\rightarrow\mathcal{Y}$ is trained to inform the choice for the next iteration.
The ultimate goal of deep active learning is to obtain high predictive accuracy for the trained neural network while annotating as few examples as possible.
@@ 3.2 Review of Imbalanced Active Learning: GALAXY
We first consider a binary case of imbalanced classification, where we assume $N_{1}<N_{2}$ without loss of generality. After a total of $T$ rounds and sampling $TB$ examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$. On the other hand, as shown previously in Zhang et al. (2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low. The reasoning is simply demonstrated in Figure 0(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples (all) from the majority class. It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
@@
Therefore, the key objective in imbalanced active learning is labeling examples that are both uncertain and class balanced. The proposed algorithm GALAXY (Zhang et al., 2022) follows the following two-phased procedure: 1. Identify the optimal separation threshold $\widehat{p}^{\star}$, as shown in Figure 0(a). 2. Annotate examples in the minority class closest to $\widehat{p}^{\star}$.
GALAXY draws inspiration from graph-based active learning in identifying $\widehat{p}^{\star}$. In particular, it follows a modified bisection procedure that sequentially annotate examples to search for all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 0(b)). GALAXY then annotate examples close to all such cuts. However, the algorithm suffers from two weaknesses. Firstly, a bisection procedure only allows sequential labeling, which prevents multiple annotators labeling in parallel. Secondly, the problem of identifying the optimal separation threshold is a noisy procedure, since the predictive sigmoid scores may not perfectly separate the minority class examples from the majority. GALAXY mitigates the problem by wastefully identifying and annotating around all cuts, which could lead to annotating large number of majority class examples (shown in Figure 0(b)).
@@ 4 A Robust Algorithm with Batch Labeling
In this section, we formally define optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem. We then propose a novel algorithm inspired by the agnostic active learning literature (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021).
@@ 4.1 An 1-D Reduction to Agnostic Active Learning
We start by considering the imbalanced binary classification setting mentioned in Section 3.2. When given a neural network model, we let $\widehat{p}:\mathcal{X}\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores. Here, a higher sigmoid score represents a higher confidence of the example being in class $2$. Let $q_{1},...,q_{N}$ denote the sigmoid scores of every example in the pool $X$, where $q_{i}=\widehat{p}(x_{i})$ for each $i\in[N]$. We also sort examples by the sigmoid predictive score similar to Section 3.2. Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots q_{(N)}$ be a sorted permutation of $q_{1},...,q_{N}$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels. We are now ready to define the optimal separation threshold as described in Section 3.2 below.
@@
Given the sorted predictive sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
$j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:1\leq i\leq j\}\rvert\right.$
|$\left.-\lvert\{y_{(i)}=2:1\leq i\leq j\}\rvert\right).$ (1)
In other words, $j^{\star}$ has the largest difference in the number of class $1$ examples from class $2$ examples to its left. Furthermore, ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. In particular, let =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ be the hypothesis class. Here each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ maps sigmoid scores to binary target classes and
@@
The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix Appendix A., we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (1). Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
Multi-Class Classification. To generalize the above problem formulation to multi-class classification, we follow a similar strategy to Zhang et al. (2022). Specifically, for each class $k$, one can view the problem of class-$k$ v.s. others as a binary classification problem. In other words, class $k$ can be seen as class $1$ and classes $\{1,...,k-1,k+1,...,N\}$ can be seen as class $2$. The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems. Moreover, let $\hat{p}:\mathcal{X}\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores. Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\hat{p}(x_{i})]_{k^{\prime}}-[\hat{p}(x_{i})]_{k}$ to sort the examples. We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
4.2 Algorithm
We are now ready to state our algorithm as shown in Algorithm 1. The algorithm follows a two-phased, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class.
@@ 4.2 Algorithm
     Initialize budget per class $b\leftarrow\frac{B_{\text{train}}}{2K}$.
@@
        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$.
        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{(x_{(i)},y_{(i)})\}_{i=1}^{N})$.
@@
     Compute budget per class $b\leftarrow\frac{B_{\text{train}}-|L_{t}|}{K}$.
@@
To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2). Specifically, we keep a version space of the hypotheses with higher likelihood of being the optimal separation threshold. We annotate examples sequentially in batches and shrink the version space by a fixed rate each iteration. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left.
To address batch labeling, we note the number of examples one collect before retraining is usually far greater than the number of annotators annotating in parallel. In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained. In practice, this number is usually determined by the practical constraints of computational training cost. On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel. As discussed above, $B_{\text{parallel}}<<B_{\text{train}}$ in practice. Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{(x_{(i)},y_{(i)})\}_{i=1}^{N}$ (ground truth labels are hidden to the learner).
  Initialize: Version space: shortest segment of indices $[I,J]$, where $X_{I},X_{J}\in L$, $\forall i\leq I,x_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,x_{(j)}\in L\implies y_{(j)}\neq k$.
@@
     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $x_{(I)}$, …, $x_{(J)}$ for annotation and insert to $L$.
@@
^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ x_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ x_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$

### 01 Feb 2024 v2

@@ Abstract
Class imbalance is a prevalent issue in real world machine learning applications, often leading to poor performance in rare and minority classes. With an abundance of wild unlabeled data, active learning is perhaps the most effective technique in solving the problem at its root – collecting a more balanced and informative set of labeled examples during annotation. Label noise is another common issue in data annotation jobs, which is especially challenging for active learning methods. In this work, we conduct the first study of active learning under both class imbalance and label noise. We propose a novel algorithm that robustly identifies the class separation threshold and annotates the most uncertain examples that are closest from it. Through a novel reduction to one-dimensional active learning, our algorithm DIRECT is able to leverage the classic active learning literature to address issues such as batch labeling and tolerance towards label noise. We present extensive experiments on imbalanced datasets with and without label noise. Our results demonstrate that DIRECT can save more than 60% of the annotation budget compared to state-of-art active learning algorithms and more than 80% of annotation budget compared to random sampling.
@@ 1 Introduction
Large-scale deep learning models are playing increasingly important roles across many industries. Human feedback and annotations have played a significant role in developing such systems through. Progressively over time, we believe the role of humans in a machine learning pipeline would shift to annotating rare yet important cases. However, under data imbalance, the typical strategy of randomly choosing examples for annotation becomes especially inefficient. This is because the majority of the labeling budget would be spent on common and well-learned classes, resulting in insufficient rare class examples for training an effective model. For many annotation jobs, this challenge of data imbalance is further compounded by label noises induced from annotators. In this paper, we mitigate these issues by proposing a deep active learning strategy. Our algorithm DIRECT sequentially and adaptively chooses informative and more class-balanced examples for annotation while being robust to noisy annotations. To the best of our knowledge, this is the first deep active learning study to address the challenging yet prevalent scenario where both imbalance and label noise coexist.
Inspired by previous work by Zhang et al. (2022), we propose a novel reduction of the imbalanced classification problem into one-dimensional active learning problems. For each class, our reduction sorts unlabeled examples into an list ordered by one-vs-rest margin scores. The goal of DIRECT is to find the optimal separation threshold which best separates the examples in any given class from the rests. With an equivalence to finding the best threshold classifier, we were able to employ classic active learning literature for obtaining the separation threshold. Once the optimal separation threshold is identified, DIRECT annotates a set of more class-balanced examples near this threshold. As examples around the optimal separation threshold tend to have high uncertainty, DIRECT annotates examples that are both class-balanced and informative.
With the reduction to classic one-dimensional active learning literature, DIRECT also inherits the advantages from classic active learning algorithms such as robustness against label noise and batch labeling. This is in contrast to the state-of-art imbalanced active learning algorithm GALAXY proposed by Zhang et al. (2022), which requires sequential labeling and is much less robust against label noise. As demonstrated in Figure 1, under both noisy and noiseless label settings, DIRECT improves significantly over GALAXY’s label efficiency – DIRECT requires less annotations to achieve the same accuracy. It achieves state-of-art label efficiency gain on both supervised fine-tuning of ResNet-18 and semi-supervised fine-tuning of large pretrained model under the LabelBench framework (Zhang et al., 2023a). Lastly, DIRECT also allows for parallel annotations with various batch sizes.
(a) Imbalanced CIFAR-10, two classes, no label noise.
(b) Imbalanced CIFAR-100, two classes, 20% label noise.
(c) LabelBench FMoW (62 classes with imbalance), no label noise
Figure 1: Performance of DIRECT over existing baselines for both noiseless and noisy settings. The x-axis represents the total number of labeled examples so far and the y-axis shows the neural network’s balanced accuracy. Both (a) and (b) are using supervised training of ResNet-18. In (c), we finetune CLIP ViT-B32 model in combination of semi-supervised training under the LabelBench framework (Zhang et al., 2023a). $B_{\text{parallel}}$ is the batch size indicating the number of parallel annotators for the annotation job. $B_{\text{parallel}}=1$ indicates the synchronous annotation requirement by GALAXY. Our algorithm DIRECT takes pre-specified $B_{\text{parallel}}$ as input, which is determined by real world decisions.
@@
  • We propose a novel reduction that bridges the advancement in theoretical active learning literature to imbalanced active classification for deep neural networks.
  • Our algorithm DIRECT addresses the prevalent imbalance and label noise issues and annotates a more class-balanced and informative set of examples.
  • Compared to state-of-art algorithm GALAXY (Zhang et al., 2022), DIRECT allows parallel annotation by multiple annotators while still maintaining significant label-efficiency improvement.
  • We conduct experiments across eight dataset settings, four levels of label noise and for both ResNet-18 and large pretrained model (CLIP ViT-B32). DIRECT consistently performs the best compared against existing baseline algorithms. Quantitatively, DIRECT can save more than $60\%$ annotation cost compared to GALAXY, and more than $80\%$ annotation cost compared to random sampling.
@@ 2 Related Work
Class-Balanced Deep Active Learning Active Learning strategies sequentially and adaptively choose examples for annotation (Settles, 2009). Many uncertainty-based deep active learning methods extend the traditional active learning literature such as margin, least confidence and entropy sampling (Tong & Koller, 2001; Settles, 2009; Balcan et al., 2006; Kremer et al., 2014). These methods have also been shown to perform among top methods when fine-tuning large pretrained models and combined with semi-supervised learning algorithms (Zhang et al., 2023a). More sophisticated methods have also been proposed to optimize the chosen unlabeled examples’ uncertainty (Gal et al., 2017; Ducoffe & Precioso, 2018; Beluch et al., 2018), diversity (Sener & Savarese, 2017; Geifman & El-Yaniv, 2017; Citovsky et al., 2021), or a mix of both (Ash et al., 2019, 2021; Wang et al., 2021; Elenter et al., 2022; Mohamadi et al., 2022). However, these methods often perform poorly under prevalent and realistic scenarios such as under label noises (Khosla et al., 2022) or under class imbalance (Kothawade et al., 2021; Zhang et al., 2022, 2023a).
@@
Deep Active Learning under Label Noise Label noisy settings has rarely been studied in the deep active learning literature. Related but tangential to our work, several papers have studied to use active learning for cleaning existing noisy labels (Lin et al., 2016; Younesian et al., 2021). In this line of work, they assume access to an oracle annotator that will provide clean labels when queried upon. This is fundamentally different from our work, where our annotator may provide noisy labels. Another line of more theoretical active learning research studies active learning with multiple annotators with different qualities (Zhang & Chaudhuri, 2015; Chen et al., 2022). The primary goal in these work is to identify examples a weak annotator and a strong annotator may disagree, in order to only use the strong annotator on those instances. In our work, we assume access to a single source of annotator that is noisy, which is prevalent in annotation jobs today. Recently, Khosla et al. (2022) proposed a novel deep active learning algorithm specialized for Heteroskedastic noise, where different “regions” of examples are subject to different levels of noise. Unlike their work, our work is agnostic to the noise distributions and conduct experiments on uniformly random corrupted labels.
To the best of our knowledge, no existing deep active learning work has studied the scenario where both imbalance and label noise present. Yet, this setting is the most prevalent in real-world annotation applications.
Agnostic Active Learning A critical part of our solution is the reduction to a one dimensional agnostic active learning problem, e.g. Section 4.1. There is extensive work on agnostic learning in general, and we refer the interested reader to the survey (Hanneke et al., 2014) for a thorough discussion. All of these works, beginning with the seminal works (Balcan et al., 2006; Dasgupta et al., 2007) follow, a familiar paradigm of disagreement based learning. This involves maintainin a version space of promising hypothesis at each time, and constructing the disagreement region as a set of unlabeled examples for which there exists two hypotheses that disagree on their predictions. The example to label is chosen by sampling from a informative sampling distribution computed over the disagreement region. Several approaches have been proposed for computing the sampling distribution, e.g. Jain & Jamieson (2019); Katz-Samuels et al. (2020, 2021); Huang et al. (2015). As described in Section 4.2, our main subroutine VReduce is equivalent to fixed-budget one dimensional threshold disagreement learning based on the ACED algorithm of Katz-Samuels et al. (2021). We remark that the problem of class imbalance has been largely ignored by the classical work on traditional agnostic learning, and this paper hopes to close this gap.
@@ 3.1 Notations
We study the pool-based active learning problem, where an initial unlabeled set of $N$ examples $X=\{x_{1},...,x_{N}\}$ are available for annotation. Their corresponding labels $Y=\{y_{1},...,y_{N}\}$ are initially unknown. Furthermore, we study the multi-class classification problem, where the space of labels $\mathcal{Y}:=[K]$ is consisted $K$ classes. Moreover, let $N_{1},...,N_{K}$ denote the number of examples in $X$ of each class. We define the imbalance ratio as $\gamma=\frac{\min_{k\in[K]}N_{k}}{\max_{k^{\prime}\in[K]}N_{k^{\prime}}}$.
A deep active learning algorithm iteratively chooses batches of examples for annotation. During the $t$-th iteration, the algorithm is given labeled and unlabeled sets of examples, $L_{t}$ and $U_{t}$ respectively, where $L_{t}\cup U_{t}=X$ and $L_{t}\cap U_{t}=\emptyset$. The algorithm then chooses $B$ examples from the unlabeled set $X^{(t)}\subseteq U_{t}$ and then obtains their corresponding labels $Y^{(t)}$. The labeled and unlabeled sets are then updated, i.e., $L_{t+1}\leftarrow L_{t}\cup X^{(t)}$ and $U_{t+1}\leftarrow U_{t}\backslash X^{(t)}$. Based on new labeled set $L_{t+1}$ and its corresponding labels, a neural network $f_{t}:X\rightarrow[K]$ is trained to inform the choice for the next iteration. The ultimate goal of deep active learning is to obtain high predictive accuracy for the trained neural network while annotating as few examples as possible.
@@ 3.2 Review of Imbalanced Active Learning: GALAXY
We first consider a binary case of imbalanced classification, where we assume $N_{1}<N_{2}$ without loss of generality. After a total of $T$ rounds and sampling $TB$ examples, random sampling would annotate a subset of $X$ with an imbalance ratio close to $\frac{N_{1}}{N_{2}}$. On the other hand, as shown previously in Zhang et al. (2022), uncertainty based methods such as confidence (Settles, 2009), margin (Tong & Koller, 2001; Balcan et al., 2006) and entropy (Kremer et al., 2014) sampling could annotate a set of examples with imbalance ratio arbitrarily low. The reasoning is simply demonstrated in Figure 1(a), where choosing examples with predictive sigmoid score $\widehat{p}$ closest to .5.5.5.5 may end up annotating examples almost all from the majority class. It is also empirically shown that annotating examples most certain to be in the minority class is not the most effective in improving model performance.
@@
Therefore, the key objective in imbalanced active learning is to label examples that are both uncertain and class balanced. In both GALAXY (Zhang et al., 2022) and this work, we decompose the objective into the following two-phased procedure: 1. Identify the optimal separation threshold $j^{\star}$ that best separates the minority class from the majority class, as shown in Figure 1(a). 2. Annotate equal number of examples next to $j^{\star}$ from both sides.
GALAXY draws inspiration from graph-based active learning in identifying $p^{\star}$. In particular, it follows a modified bisection procedure that sequentially annotates examples to find all cuts, namely thresholds with a class 1 example to the left and a class 2 example to the right (see Figure 1(b)). As the optimal separation threshold must be located at one of these cuts, GALAXY annotates examples close to all the found cuts. However, the algorithm suffers from three weaknesses. First, as shown in Figure 1(b), by annotating around all cuts, the algorithm may waste a significant portion of the annotation budget around outliers, which could lead to annotating large number of majority class examples. Second, the above problem can be magnified when annotators provide noisy labels. Third, a bisection procedure only allows for sequential labeling, which prevents multiple annotators labeling in parallel. In this paper, we take a DIRECT approach by identifying only the optimal separation threshold and address all of the shortcomings above.
@@ 4 A Robust Algorithm with Batch Labeling
In this section, we formally define the optimal separation threshold and pose the problem of identifying it as an 1-dimensional reduction to the agnostic active learning problem. We then propose an algorithm inspired by the agnostic active learning literature (Balcan et al., 2006; Dasgupta et al., 2007; Hanneke et al., 2014; Katz-Samuels et al., 2021).
@@ 4.1 An 1-D Reduction to Agnostic Active Learning
We start by considering the imbalanced binary classification setting mentioned in Section 3.2. When given a neural network model, we let $\widehat{p}:X\rightarrow[0,1]$ be the predictive function mapping examples to sigmoid scores. Here, a higher sigmoid score represents a higher confidence of the example being in class $2$. We sort examples by their sigmoid predictive score similar to Section 3.2. Formally, we let $q_{(1)}\leq q_{(2)}\leq\cdots\leq q_{(N)}$ be a sorted permutation of $\widehat{p}(x_{1}),...,\widehat{p}(x_{N})$, and let $y_{(1)},y_{(2)},...,y_{(N)}$ denote the sorted list’s corresponding labels. We are now ready to define the optimal separation threshold as described in Section 3.2.
@@
Given a sorting of $X$ the points according the sigmoid scores $0\leq q_{(1)}\leq\cdots\leq q_{(N)}\leq 1$, we define the optimal separation threshold as $j^{\star}\in[N]$ such that
$j^{\star}=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=1:i\leq j\}\rvert-\lvert\{y_{(i)}=2:i\leq j\}\rvert\right)$
$=\operatorname*{arg\,max}_{j}$ (|$\left(\lvert\{y_{(i)}=2:i>j\}\rvert-\lvert\{y_{(i)}=2:i>j\}\rvert\right).$ (1)
In other words, on either side of $j^{\ast}$, it has the largest discrepancy in the number of examples between the two classes. This captures the intuition of Figure 1(b) — our goal is to find a threshold that best separates one class from the other. We quickly remark that ties are broken by choosing the largest $j^{\star}$ that attains the argmax if class $1$ is the minority class and the lowest $j^{\star}$ otherwise.
1D Reduction. We now provide a reduction of finding $j^{\star}$ to an 1-dimensional agnostic active learning problem. We define the hypothesis class =$\mathcal{H}=\{h_{0},h_{1},...,h_{N}\}$ where each hypothesis $h_{j}:[0,1]\rightarrow\{1,2\}$ is defined as
@@
and $q_{(0)}=0$ defines the hypothesis $h_{0}$ that predicts class $2$ at all times. The empirical zero-one loss for each hypothesis is then defined as $\mathcal{L}(h_{j})=\sum_{i=1}^{N}\mathbf{1}\{h_{j}(q_{(i)})\neq y_{(i)}\}$. In Appendix A, we show that optimizing for the zero-one loss $\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$ is equivalent to (4.1). Namely, $j^{\star}\in\operatorname*{arg\,min}_{0\leq j\leq N}\mathcal{L}(h_{j})$, transforming the problem of finding the optimal separation threshold to an 1-D agnostic active learning problem.
Multi-Class Classification. To generalize the above problem formulation to multi-class classification, we follow a similar strategy to Zhang et al. (2022). Specifically, for each class $k$, one can view the problem of class-$k$ v.s. others as a binary classification problem. In other words, class $k$ can be seen as class $1$ and classes $\{1,...,k-1,k+1,...,N\}$ can be seen as class $2$. The goal therefore becomes finding all $K$ optimal separation thresholds, which is equivalent with solving $K$ 1-D agnostic active learning problems. Moreover, let $\widehat{p}:X\rightarrow\Delta^{(K-1)}$ denote the neural network prediction function mapping examples to sigmoid scores. Instead of using sigmoid scores, for class $k$, we use the margin scores $q_{i}^{k}:=\max_{k^{\prime}}[\widehat{p}(x_{i})]_{k^{\prime}}-[\widehat{p}(x_{i})]_{k}$ to sort the examples. When multiple examples have the same margin score, we further break ties by their corresponding confidence scores $[\widehat{p}(x_{i})]_{k}$. We also note that sorting with margin scores is equivalent to sorting with sigmoid scores in the binary classification setting, so this is a strictly more general formulation to multi-class classification.
@@
     Initialize budget per class $b\leftarrow B_{\text{train}}/2K$.
@@
        Compute margin scores $q_{1}^{k},...,q_{N}^{k}$ based on $f_{t-1}$ and sort them by $q_{(1)}^{k}\leq\cdots\leq q_{(N)}^{k}$ and break ties so that $q_{(i)}^{k}=q_{(i+1)}^{k}\implies[\widehat{p}(x_{(i)})]_{k}\geq[\widehat{p}(x_{(i+1)})]_{k}$.
        Find optimal separation threshold: $L_{t}\leftarrow\text{VReduce}(L_{t},b,k,B_{\text{parallel}},\{q^{k}_{(i)}\}_{i=1}^{N})$.
@@
     Compute budget per class $b\leftarrow(B_{\text{train}}-|L_{t}|)/K$.
@@
  Input: Labeled set $L$, budget $b$, class of interest $k$, parallel batch size $B_{\text{parallel}}$, examples and ground truth labels sorted by uncertainty $\{q_{(i)}\}_{i=1}^{N}$.
  Define: For examples in the labeled set, we let $y_{(i)}$ denote the label of the example corresponding to $q_{(i)}$.
  Initialize: Version space: shortest segment of indices $[I,J]$, with $q_{(I)},q_{(J)}$ $\forall i\leq I,q_{(i)}\in L\implies y_{(i)}=k$ and $\forall j\geq J,q_{(j)}\in L\implies y_{(j)}\neq k$.
@@
     Sample uniformly at random $B_{\text{parallel}}$ unlabeled examples in $q_{(I)}$, …, $q_{(J)}$ for annotation and insert to $L$.
@@
^$\widehat{\mathcal{L}}^{k}(s)=\sum_{\begin{subarray}{c}r\leq s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}\neq k\}+\sum_{\begin{subarray}{c}r>s:\\ q_{(r)}\in L\end{subarray}}\mathbf{1}\{y_{(r)}=k\}.$
@@
4.2 Algorithm
We are now ready to state our algorithm as shown in Algorithm 1. The algorithm follows a two-phased procedure, where the first phase aims to identify the optimal separation threshold for each class. The second phase then annotates examples closest to the estimated optimal separation thresholds for each class. Our set of labeled examples at anytime are maintained in $L_{t}$.
To identify the optimal separation threshold, we loop over each class $k$ and run a fix-budget batch active learning procedure for the one-dimensional reduction (as shown in Algorithm 2) on the set of labeled examples. Our method is motivated by the fixed-budget algorithm for active classification, ACED (Katz-Samuels et al., 2021). Intuitively, in the non-noisy or separable setting, where the loss of $\mathcal{L}(h_{j^{\star}})=0$, we could run a binary search procedure to learn $j^{\star}$. However, we can’t assume that $j^{\star}$ perfectly separates the classes. ACED circumvents this by maintaining a version space of threshold classifiers, namely an interval $[I,J]$. In each of the $b/B_{\text{parallel}}$ rounds, $B_{\text{parallel}}$ samples are taken and appended to the set of labeled examples $L$. Then, an interval $I^{\prime},J^{\prime}$ is computed, whose length is a factor of $1/c$ smaller than the current $I,J$ of low-loss classifiers. The shrinkage rate $c$ is determined by the budget and batch size, so that after the final iteration, the version space has exactly one hypothesis left. The fixed confidence (or alternatively $\delta$-PAC) analogue of ACED is known to match instance optimal lower bounds for active classification up to logarithmic factors. Under benign noise conditions, the probability of misidentifying the optimal separation threshold decays exponentially w.r.t. budget $b$.
To address batch labeling, we note the number of examples collected before retraining is usually far greater than the number of annotators annotating in parallel. In particular, we let $B_{\text{train}}$ denote the number of examples the algorithm collects before the neural network is retrained. In practice, this number is usually determined by the constraints of computational training cost. On the other hand, we let $B_{\text{parallel}}$ denote the number of examples annotated in parallel. As discussed above, $B_{\text{parallel}}\ll B_{\text{train}}$ in practice. Lastly, as will be discussed in Section 6, a variant of our algorithm can also be proposed for asynchronous labeling.